# Improving Nonpreemptive Multiserver Job Scheduling with Quickswap

### v1

@@ 7 Conclusion
Note that $H_{3,k}\sim B^{S}_{S_{1}}$ as the time going from having $k$ to $k-1$ small jobs is the busy period because the completion rate of small jobs here is $n\mu_{1}$. Then, the phase duration of phase 3 is the sum of these periods. Formally, $H_{3}=\sum\nolimits_{j=\ell+1}^{n-1}H_{3,j}.$ Then, by standard transform techniques, Section 5.4 holds. ∎
@@
In phase 4, no further small job arrivals are allowed into service, and there are $\ell$ small jobs at the beginning of phase 4. Therefore, we can write $H_{4}$ as a sum of i.i.d. exponential distributions: $H_{4}=\sum_{j=1}^{\ell}Exp(j\mu_{1}).$ Therefore, by standard transform techniques, $\widetilde{H_{4}}(s)=\prod_{j=1}^{\ell}\widetilde{Exp(j\mu_{1})}(s)=\prod_{j=1}^{\ell}\frac{j\mu_{1}}{j\mu_{1}+s}.$

### v2

@@ Acknowledgements
Note that $H_{3,k}\sim B^{L}_{S_{1}}$ as the time going from having $k$ to $k-1$ light jobs is the busy period because the completion rate of light jobs here is $k\mu_{1}$. Then, the phase duration of phase 3 is the sum of these periods. Formally, $H_{3}=\sum\nolimits_{j=\ell+1}^{k-1}H_{3,j}.$ Then, by standard transform techniques, Section 5.4 holds. ∎
@@
In phase 4, no further light job arrivals are allowed into service, and there are $\ell$ light jobs at the beginning of phase 4. Therefore, we can write $H_{4}$ as a sum of i.i.d. exponential distributions: $H_{4}=\sum_{j=1}^{\ell}Exp(j\mu_{1}).$ Therefore, by standard transform techniques, $\widetilde{H_{4}}(s)=\prod_{j=1}^{\ell}\widetilde{Exp(j\mu_{1})}(s)=\prod_{j=1}^{\ell}\frac{j\mu_{1}}{j\mu_{1}+s}.$
Appendix C Fairness Evaluation
To compare the fairness of scheduling policies, we compute Jain’s Fairness index [jain1998quantitativemeasurefairnessdiscrimination] as
$J(\mathbb{E}[T^{(1)}],\mathbb{E}[T^{(2)}],\cdots,\mathbb{E}[T^{(k)}])=\frac{\left(\sum_{j=1}^{k}\mathbb{E}[T^{(j)}]\right)^{2}}{k\sum_{j=1}^{k}(\mathbb{E}[T^{(j)}])^{2}}.$ (8)
The value of the fairness index is between $\frac{1}{k}$ and 1. A higher value of the fairness index indicates that the scheduling policy is fairer.
We examine the fairness of various scheduling policies in Figure 7. Although MSF and First-Fit achieve low unweighted mean response time (Figure 7(a)), the heavy jobs experience orders of magnitude larger waiting times than the light jobs (Figure 7(b)) under these policies. Under Adaptive and Static Quickswap, on the other hand, the mean response times of light and heavy jobs are comparable. As a result, Adaptive and Static Quickswap achieve higher fairness indices compared to MSF and First-Fit (Figure 7(c)). Because unweighted mean response time can hide these important imbalances between the job classes, our evaluation uses weighted mean response time — a metric that balances overall mean response time and fairness.
(a) Unweighted overall mean response time
(b) Mean response time by classes
(c) Jain’s Fairness Index
Figure 7: The response time performance and fairness index as a function of the overall arrival rate for MSJ systems serving a Google Borg workload. The middle plot, Figure 7(b), shows the mean response time by class, where the dotted lines denote the mean response time of the heaviest jobs and the solid lines denote the mean response time of the lightest jobs. The other plots, left and right, show combined mean response time and fairness metrics across classes in a single, solid line.
Appendix D Comparison with Preemptive Policies
Although preemption is either infeasible or carries significant overhead for many datacenter workloads, we compare Adaptive and Static Quickswap to a preemptive policy for the sake of completeness. Here, we assume that the preemptive policy, ServerFilling, can preempt jobs with no overhead or setup cost. We see that ServerFilling can use preemption to greatly outperform any non-preemptive policy.
(a) Unweighted mean response time
(b) Weighted mean response time
Figure 8: The overall mean response time of the system as a function of the overall arrival rate for MSJ systems serving a Google Borg workload. Points of higher arrival rates for the First-Fit policy are hidden because the experiments did not converge.
Figure 8 shows that the preemptive ServerFilling policy [wcfs] greatly outperforms all non-preemptive policies with respect to both unweighted and weighted mean response time. ServerFilling uses preemptions to guarantee full resource utilization whenever there are more than $k$ jobs in the system. The non-preemptive scheduling policies, on the other hand, may waste significant service capacity even when there are many jobs in the queue. This waste leads to higher mean response times for all non-preemptive policies, including our Quickswap policies.